DQuestion 1 1 pts 7.1 sound is the common name for an eight-channel surround audio system commonly used in home theater configurations. It adds two additional speakers to the more conventional six-channel (5.1) audio configuration. As with 5.1 surround sound, 7.1 surround sound positional audio uses the standard front, center, and LFE (subwoofer) speaker configuration. However, whereas a 5.1 surround sound system combines both surround and rear channel effects into two channels (commonly configured in home theatre set-ups as two rear surround speakers), a 7.1 surround system splits the surround and rear channel information into four distinct channels, in which sound effects are directed to left and right surround channels, plus two rear surround channels. In Blu-ray Disc and HD DVD home video formats, a lossless 7.1 PCM audio is available. Audio consists of 8 multiplexed identical channels each with 48 ksamples/s. The overall bit rate is 6.144 Mbits/s. How many quantization levels are there on each channel?
media%2F467%2F467262bb-603f-4d57-9d69-64
Answer #1

each channel is having a sampling rate 48kS/s and bit rate 6.144Mb/s.

therefore required resolution is n- bits = bit-rate/sampling-rate=6.144M/48k = 128 bits

the no. of quantization levels = 2^n = 2^128 = 3.4028236692093846346337460743177e+38

* answer not given in the options provided